# Helical Ribbons: Simple Chiral Sedimentation

### v1

@@ Abstract
There are helical ribbons with special lengths that have axisymmetric translation-rotation coupling. At these lengths, a bifurcation occurs where the fixed points in tilt-spin space switch between centers and saddles. These axisymmetric helicoids have a particularly simple $\mathbb{b}$ tensor and can function as a first step on the roadmap that extends to general helical ribbons with triaxial $\mathbb{b}$ and on to yet more general shapes that are not co-centered. We also find that helical ribbons with $b_{11}$ and $b_{22}$ of opposite signs can have trajectories whose horizontal center-of-mass displacement is unbounded.
Experiments that precisely fabricate complex shapes and measure full mobility tensors have now revealed several unexpected features of translation-rotation coupling during sedimentation Collins et al. (2021); Miara et al. (2024). Together with the present work, these point to an important design challenge to create geometries that optimize translation-rotation coupling. Some progress in this direction has been made Keaveny et al. (2013), but the challenge goes far beyond optimizing propulsion efficiency in one dimension. For co-centered particles, each eigenvalue of $\mathbb{b}$ can be optimized relative to the others to achieve different goals. A full picture of which geometries optimize translation-rotation coupling in Stokes flow could be a major step toward the broader goal of developing better ways to quantify chiral geometry.
I Supplemental Material
I.1 Reconstruction of 3D Trajectories
We use three Phantom VEO 640 cameras with nearly orthogonal viewing angles to capture 3D dynamics. LED panels provide bright field illumination onto which black helical ribbons produce clean images. We use methods of camera calibration and orientation measurement developed in Oehmke et al. (2021).
Initial camera calibration parameters are obtained using a static two-plane calibration and then they are refined to sub-pixel accuracy with a dynamic calibration using images of small settling metal spheres.
To find the position and orientation of the helical ribbon from video images, we project a set of points on the surface of an ideal 3D helical ribbon onto the image planes of the three cameras using the camera calibration. The average distance of each these points from a bright pixel in the thresholded image is minimized to find the 3 position and 3 orientation coordinates.
I.2 Fabricating Co-Centered Particles
We tried three different methods for fabricating helical ribbons. The first was mechanical twisting of polystyrene strips. Strips were cut on a paper cutter, twisted with a mechanical jig, annealed at 90o C, and then cut to the desired length. This produced cost effective helical ribbons with variability in width, length, and pitch that required optical sorting to select specific particles. There were also problems with non-reproducible sedimentation dynamics, likely a result of the thin polystyrene having mobility tensors that are very sensitive to tiny imperfections, lint, or bubbles.
@@ I.2 Fabricating Co-Centered Particles
I.3 Mobility tensor normalization
We measure diagonal elements of the $\mathbb{a}$ and $\mathbb{b}$ mobility tensors and normalize them by tr($\mathbb{a}$)/3 and $b_{33}$ respectively. Physically, tr($\mathbb{a}$)/3 is the average drag over all orientations. However, tr($\mathbb{b}$) can be zero, so this is not a good normalization for $\mathbb{b}$. Instead, for this particle we choose to normalize by $b_{33}$, because it is larger than the other two eigenvalues and its sign is preserved as we change the helical ribbon length. In the limit of infinite length, $b_{33}$ is the only non-zero eigenvalue.
Figure 6: Normalization factors for mobility tensors as a function of helical ribbon length. (a) Average of the eigenvalues of $\mathbb{a}$ (b) Largest eigenvalue of $\mathbb{b}$.
Figure 6 provides absolute scaling values of the normalization factors used for $\mathbb{a}$ and $\mathbb{b}$. Notably, the helical ribbon’s mass increases faster than its drag as length increases, resulting in smaller sedimentation velocities for shorter particles.
I.4 Movies of spatial trajectories
The spatial trajectories of an arbitrary particle fall in a hierarchy of complexity governed by the number of distinct $\mathbb{a}$ and $\mathbb{b}$ eigenvalues. $\mathbb{b}$ determines what frequencies are present and $\mathbb{a}$ determines if those frequencies can couple to spatial dynamics.
To communicate the wide range of spatial trajectories that are possible, we created movies showing how the trajectories change as the initial orientation is continuously changed. Movies are available at: https://www.youtube.com/channel/UCVRg1jeDOSidH7oYnLdqZ2Q. The movies show simulated trajectories using mobility tensors measured with fits to experimental data for each of the four helical ribbon lengths that we measured. The shortest and longest particles show switching of the handedness of spatial trajectories because the values of $b_{11}$ and $b_{22}$ (shown in Fig. 3 in the main article) have opposite signs. This allows cancellation of contributions to the change in $\phi$ along the trajectory. The other two particles always have left handed trajectories because the negative eigenvalues dominate.
Additionally, a video of the example trajectory shown in Fig. 1 is available at: https://www.youtube.com/watch?v=n9FWTLHiUK4. The angular phase space is pictured at the bottom left with the same axes as in Fig. 1(d).
Acknowledgements.
We thank Bennet Grossman for assistance in setting up the experimental apparatus and Gleb Shevchuk for conversations about advanced 3D printing techniques. We acknowledge support from the NSF under grant DMR-1508575 and CBET-2211704, the Army Research Office under grant W911NF-17-1-0176, and Vetenskapsrådet under grant no. 2021-4452.

### v2

@@ IV Conclusion
There are helical ribbons with special lengths that have axisymmetric translation-rotation coupling. At these lengths, a bifurcation occurs where the fixed points in tilt-spin space switch between centers and saddles. These axisymmetric helicoids have a particularly simple $\mathbb{b}$ tensor and can function as another step on the roadmap that extends to general helical ribbons with triaxial $\mathbb{b}$ and on to yet more general shapes that are not co-centered.
Experiments that precisely fabricate complex shapes and measure full mobility tensors have now revealed several unexpected features of translation-rotation coupling during sedimentation Collins et al. (2021); Miara et al. (2024). Together with the present work, these point to an important design challenge to create geometries that optimize translation-rotation coupling. Some progress in this direction has been made Keaveny et al. (2013), but the challenge goes far beyond optimizing propulsion efficiency in one dimension. For co-centered particles, each eigenvalue of $\mathbb{b}$ can be optimized relative to the others to achieve different goals. A full picture of which geometries optimize translation-rotation coupling in Stokes flow could be a major step toward the broader goal of developing better ways to quantify and design chiral geometries.
Acknowledgements.
We thank Bennet Grossman for assistance in setting up the experimental apparatus, Gleb Shevchuk for conversations about advanced 3D printing techniques and G. Sarnitsky for discussions. We acknowledge support from the NSF under grant DMR-1508575 and CBET-2211704, the Army Research Office under grant W911NF-17-1-0176, the Initiative d’Excellence d’Aix-Marseille Université - A*MIDEX - AMX-19-IET-010, the French National Research Agency (ANR) in the framework of NetFlex Project (ANR-21-CE30-0040) and Vetenskapsrådet under grant no. 2021-4452.
Appendix A Fabricating Co-Centered Particles
We tried three different methods for fabricating helical ribbons. The first was mechanical twisting of polystyrene strips. Strips of polystyrene sheet were cut on a paper cutter, twisted with a mechanical jig, annealed at 90o C, and then cut to the desired length. This produced cost effective helical ribbons with variability in width, length, and pitch that required optical sorting to select specific particles. There were also problems with non-reproducible sedimentation dynamics, likely a result of the thin polystyrene having mobility tensors that are very sensitive to tiny imperfections, lint, or bubbles.